# Retinex-MEF: Retinex-based Glare Effects Aware Unsupervised Multi-Exposure Image Fusion

### v1

@@ Abstract
Multi-exposure image fusion consolidates multiple low dynamic range images of the same scene into a singular high dynamic range image. Retinex theory, which separates image illumination from scene reflectance, is naturally adopted to ensure consistent scene representation and effective information fusion across varied exposure levels. However, the conventional pixel-wise multiplication of illumination and reflectance inadequately models the glare effect induced by overexposure. To better adapt this theory for multi-exposure image fusion, we introduce an unsupervised and controllable method termed (Retinex-MEF). Specifically, our method decomposes multi-exposure images into separate illumination components and a shared reflectance component, and effectively modeling the glare induced by overexposure. Employing a bidirectional loss constraint to learn the common reflectance component, our approach effectively mitigates the glare effect. Furthermore, we establish a controllable exposure fusion criterion, enabling global exposure adjustments while preserving contrast, thus overcoming the constraints of fixed-level fusion. A series of experiments across multiple datasets, including underexposure-overexposure fusion, exposure control fusion, and homogeneous extreme exposure fusion, demonstrate the effective decomposition and flexible fusion capability of our model. The code will be released.
@@ 1 Introduction
Current photographic equipment often captures a dynamic range far lower than natural scenes, failing to adapt to large-scale illumination variations in the real world. This results in poor visibility and detail loss [12, 41, 21]. Multi-exposure image fusion (MEIF) merges multiple images from the same scene, taken at different exposure levels, to create a single high dynamic range (HDR) image that retains details in both highlights and shadows, enhancing dynamic range and color accuracy [2, 19, 9, 31, 40, 23, 33, 22, 17, 36, 39, 45]. This technique excels under extreme lighting conditions or when the illumination changes rapidly, offering greater flexibility in capturing scenes [35, 48, 4]. Emerging deep learning-based MEIF techniques, although promising, face limitations as supervised methods depend on ground truths that are artificially selected or synthesized, restricting their effectiveness in accurately capturing scene details and adjusting to natural exposure variations [20, 26, 55, 49, 11, 50, 27]. Unsupervised methods, using feature-level or pixel-level constraints, still suffer from image degradation due to extreme exposures, failing to fully preserve scene information. Overcoming the reliance on ground truth and accurately recovering scene information under extreme exposures remains a significant challenge in MEIF.
Figure 1: Traditional Retinex-based techniques for multi-exposure image fusion struggle with overexposed regions, which can adversely affect the learning of the reflectance component $R$. Our model circumvents this issue by modeling the glare effects $G$, effectively restoring the shared reflectance component $\hat{R}$ and achieving satisfying results.
@@
Considering that multi-exposure image sequences capture the same scene at different exposure levels, we utilize Retinex theory, which simulates the human visual system’s processing mechanisms to enhance color and brightness. This theory is extensively applied in applications such as low-light enhancement, dynamic range extension, and more [47, 34, 28, 44]. Retinex theory states that an observed image $I$ is the element-wise product of the illumination component $L$ and the scene reflectance $R$, i.e., $I=L\cdot R$ [7, 18, 21]. Here, $L$ represents the scene’s lighting conditions, influenced by the exposure level of a single capture, while $R$ represents the physical properties of the scene, which remain consistent under different lighting conditions. Within the Retinex framework, multi-exposure image sequences share a common reflectance component that captures the physical properties of the scene, while the primary distinction among the images is their varying exposure levels. Retinex theory effectively decouples exposure conditions from scene information, eliminating the need for manually designed features and allowing scene details to be revealed without relying on ground truth. Moreover, by isolating the illumination component, the fusion process becomes more flexible and targeted, enabling the generation of images that are not confined to any specific exposure level. Thus, this paper addresses the unsupervised MEIF problem using Retinex imaging theory, processing $L$ and $R$ separately for more accurate and flexible fusion.
@@
Historically, Retinex theory has been predominantly applied to low-light enhancement, where the degradation primarily results from reduced illumination levels. However, in MEIF, excessive exposure can lead to glare effects, which elevate pixel values and distort the extraction of the reflectance component $R$, as demonstrated in Fig. 2. This overexposure disrupts the learning of a consistent $R$, causing color shifts and loss of detail in Fig. 1. Therefore, in this paper, we also model the glare phenomenon induced by extreme exposure to accommodate the challenges posed by overexposure in MEIF.
@@
This paper introduces an unsupervised multi-exposure image fusion method utilizing Retinex theory, tailored to address the specific challenges of multi-exposure fusion. Our framework models glare effects from excessive exposure, allowing the extracted reflectance component $R$ to align with a shared component $\hat{R}$ instead of converging [8]. To mitigate glare, we employ a bidirectional loss on $\hat{R}$ to preserve the scene’s colors and details. Furthermore, we implement a parameterized illumination fusion criterion that ensures brightness variation and fidelity, facilitating customized exposure adjustments across the image. The contributions of this paper are summarized as follows:
@@
1) We introduce an effective Retinex-based unsupervised fusion method that separates reflectance from varying exposure levels, enabling precise exposure blending and improved detail preservation.
@@
2) We propose a fusion strategy using a symmetric and monotonic function to achieve manually controllable exposure adjustments across the image while ensuring contrast is maintained.
3) We tackle the glare issue in overexposure scenarios by discarding traditional pixel-level constraints and applying bidirectional constraints to better model and restore reflectance without glare effects.
4) Our algorithm shows excellent fusion performance on public multi-exposure fusion datasets, with exposure control experiments demonstrating its controllability.
@@
Figure 3: The workflow for the proposed Retinex-MEF method. Our method includes three modules: the Illumination Estimation Module for estimating illumination, the Shared Reflection Extraction Module for extracting the common reflectance component of the scene, and the Glare Estimation Module for glare estimation. An image with a random exposure level from multi-exposure images is sampled in each step, enhancing the ability of each module to estimate different components through Retinex-based decomposition and reconstruction.
@@ 2 Related Work
2.1 Deep learning-based MEIF
In recent years, deep learning-based multi-exposure methods can be categorized into branches of discriminative models [13], generative models [50, 42], and unified fusion models [60, 1]. Discriminative models typically rely on Convolutional Neural Networks (CNNs) [16] or transformers to predict decision maps [32, 13, 32] or directly reconstruct fused images. Supervised learning methods depend on constraints from GT and prior knowledge for effective reconstruction, while unsupervised methods leverage the similarity constraints between the fused image and source images, such as MEF-SSIM [30], to perform regression [38] and achieve more robust results. Generative models integrate multi-source images at the distribution level, enforcing consistency between the fused image and source images in terms of distribution. These models are often based on Generative Adversarial Networks (GANs) [50] or diffusion denoising models [42]. Unified fusion models [5, 49, 61] also provide insights for multi-exposure image fusion. For instance, they may use the reconstruction of source images from the fused image to aid fusion [54, 1], employ self-supervised methods using masked images to recover the original image, thereby eliminating the need for GT [25], or embed text features from large language models [60] and adaptively learn fusion losses based on the fusion task [1].
@@ 2.2 Retinex theory
Retinex theory is frequently utilized to enhance low-light conditions. Single-Scale Retinex [15] employs a fixed-scale Gaussian filter to mimic human perception of brightness, improving image contrast and brightness, yet often resulting in color distortion. Multi-Scale Retinex [37] combines filters at multiple scales to better handle image details and colors. In deep learning frameworks, Retinex theory is commonly used to estimate illumination maps and reflectance [10, 57, 46, 6, 59]. Retinexformer enhances images using a Retinex-based transformer, while EnlightenGAN [14] learns the transformation of images from low to normal lighting conditions using a Generative Adversarial Network, and Diff-Retinex enhances low-light images with a generative diffusion model. Additionally, URetinex [47] explores the use of a Retinex-based deep unfolding network for single-image low-light enhancement, and CRetinex [51] introduces a new color shift-aware Retinex model that focuses on resolving color shift issues under low-light conditions. The most relevant to this paper is the method of single-image low-light enhancement based on Retinex, trained on paired low-light images to explore the common reflectance in multiple underexposed images and enhance them in an unsupervised manner [8]. In this paper, based on Retinex theory, we decouple and separately process the illumination and reflectance of source image pairs. By extracting the common reflectance $\hat{R}$ that is unaffected by overexposure and controllably fusing illumination from multiple source images, we effectively improve image quality.
@@ 2.3 Comparison with Existing Approaches
Compared to existing multi-exposure image fusion methods, our approach does not rely on pixel-level constraints of the source images. Instead, we embrace Retinex theory to adjust exposure levels and restore the scene. By modeling the common reflectance component and the glare phenomenon caused by exposure, we restore the common reflectance component $\hat{R}$ that is unaffected by exposure levels. The adjustable fusion of illumination with $\hat{R}$ leads to superior performance in image fusion. Unlike previous Retinex-based methods, we address the glare phenomenon under extreme lighting conditions and effectively model it. Through bidirectional loss constraints, we extract the common reflectance component $\hat{R}$, providing a more effective Retinex modeling approach for multi-exposure image fusion.
@@ 3 Method
Our model is capable of handling multi-level exposure inputs. However, for clarity and comparative experimental purposes, we demonstrate using dual-image fusion as an example. Besides the overexposed and underexposed image pairs $\{I_{u},I_{o}\}$, our method incorporates an auxiliary input $I_{aux}$, derived from the same sequence $\{I_{j}\}_{j=1}^{J}$ as $\{I_{u},I_{o}\}$. The decomposition and reconstruction of varying exposure levels within $I_{aux}$ enhance our method’s ability to develop a more comprehensive and holistic understanding, thereby ensuring accurate estimation of both exposure and reflectance. However, if the image sequence lacks additional exposure levels, $I_{aux}$ may be either $I_{u}$ or $I_{o}$. During the training process of our framework, $I_{aux}$ is randomly sampled from the original sequence $\{I_{j}\}_{j=1}^{J}$ and can be the same as either $I_{u}$ or $I_{o}$.
@@ 3.1 Overview
Here, $\hat{R}$ symbolizes the intrinsic information of the scene that remains constant, while $G>0$ represents the glare effect induced by exposure. Incorporating $G$ leads to an increase in pixel values in overexposed regions.
@@ 3.2 Retinex-MEF
Estimation of Illumination. Beyond the underexposed and overexposed image pairs ${I_{u},I_{o}}$, we have another auxiliary image $I_{aux}\in\{I_{j}\}_{j=1}^{J}$. The exposure level of this auxiliary image is arbitrary and does not require differentiation from $\{I_{u},I_{o}\}$, as its primary function is to facilitate the precise estimation of both illumination and reflectance. The illumination extraction network $\mathcal{N_{I}}(\cdot)$ is composed of five convolutional layers and a $Sigmoid(\cdot)$ activation function, designed to estimate the single-channel illumination $L$ of the auxiliary image $I_{aux}$:
@@ 3.3 Training
According to Eq. (2), the ground truth for $\hat{R}$ is designated as $\hat{R}_{gt}$, leading to $I=L\cdot(\hat{R}_{gt}+G)$. To align $\hat{R}$ closely with $\hat{R}_{gt}$, it is necessary to implement a constraint. Glare effect $G$ is not consistent across all exposure levels but is predominant in overexposed images. Thus, at lower exposure levels, $G$ tends towards zero. Consequently, for most instances, $L\cdot\hat{R}_{gt}\leq L\cdot(\hat{R}_{gt}+G)=I$, though occasionally, $L\cdot\hat{R}_{gt}=I$. By applying the constraint $L\cdot\hat{R}\leq I$ universally, we ensure that $\hat{R}\leq\hat{R}_{gt}$. To enforce this, we introduce a suppression loss that caps the upper limit of $\hat{R}$, defined as:
@@
The constraint delineated in Eq. (11), which is relevant across all exposure levels, mandates that $\hat{R}\leq\hat{R}_{gt}$. Furthermore, in line with the principles of Retinex decomposition, the estimated value of $R$ should be consistent with $\hat{R}$ [47, 52, 8]. Thus, we introduce a consistency loss, assigned a lower weight than $\mathcal{L}_{suppress}$, to ensure adherence to inherent scene consistency mandated by Retinex theory:
@@
This loss ensures that the perturbed reflectance $R$, extracted at varying exposure levels, remains consistent with the shared reflectance $\hat{R}$. Given that $\hat{R}_{gt}<R$, the objective of this loss is to maximize $\hat{R}$ within the bounds set by $\hat{R}_{gt}$. The constraints $\mathcal{L}_{suppress}$ and $\mathcal{L}_{consist}$ synergistically regulate $\hat{R}$, ensuring it closely approximates $\hat{R}_{gt}$. On one hand, with both constraints using the $\ell_{1}$-norm but $\mathcal{L}_{consist}$ assigned a smaller weight than $\mathcal{L}_{suppress}$, $\mathcal{L}_{suppress}$ predominates in conflicts, maintaining $\hat{R}\leq\hat{R}_{gt}$. Conversely, when $\hat{R}\leq\hat{R}_{gt}$ has been achieved and considering $\hat{R}\leq R$, $\mathcal{L}_{consist}$ acts to elevate $\hat{R}$, guiding it towards convergence with $\hat{R}_{gt}$. This synergistic effect is illustrated in Fig. 4.
Figure 4: A schematic illustration of the synergistic effect between suppression loss $\mathcal{L}_{suppress}$ and consistency loss $\mathcal{L}_{consist}$: When $\hat{R}$ exceeds $\hat{R}_{gt}$, the two losses conflict, but $\mathcal{L}_{suppress}$ plays a dominant role. Conversely, when $\hat{R}$ is less than $\hat{R}_{gt}$, $\mathcal{L}_{consist}$ ensures that $\hat{R}$ increases.
@@
where $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}$ are the weighting coefficients, with $\alpha_{3}>\alpha_{4}$.
@@
MEFB [56] Multi-Exposure Image Fusion Dataset SICE [3] Multi-Exposure Image Fusion Dataset

### v2

@@ Abstract
Multi-exposure image fusion (MEF) synthesizes multiple, differently exposed images of the same scene into a single, well-exposed composite. Retinex theory, which separates image illumination from scene reflectance, provides a natural framework to ensure consistent scene representation and effective information fusion across varied exposure levels. However, the conventional pixel-wise multiplication of illumination and reflectance inadequately models the glare effect induced by overexposure. To address this limitation, we introduce an unsupervised and controllable method termed Retinex-MEF. Specifically, our method decomposes multi-exposure images into separate illumination components with a shared reflectance component, and effectively models the glare induced by overexposure. The shared reflectance is learned via a bidirectional loss, which enables our approach to effectively mitigate the glare effect. Furthermore, we introduce a controllable exposure fusion criterion, enabling global exposure adjustments while preserving contrast, thus overcoming the constraints of a fixed exposure level. Extensive experiments on diverse datasets, including underexposure-overexposure fusion, exposure controlled fusion, and homogeneous extreme exposure fusion, demonstrate the effective decomposition and flexible fusion capability of our model. The code is available at https://github.com/HaowenBai/Retinex-MEF.
@@ 1 Introduction
Conventional imaging sensors possess a limited dynamic range, often failing to capture scenes with significant illumination variations, resulting in detail loss in highlights or shadows [79, 17, 51, 26, 77]. Multi-exposure image fusion mitigates this by merging a sequence of images captured at different exposure levels to synthesize a single, high-quality composite [4, 24, 13, 43, 50, 32, 45, 27, 22, 47, 7, 58, 56]. The fused image thereby retains crucial details from both highlight and shadow regions, making it a vital technique for capturing scenes under challenging lighting conditions [60, 8, 41, 38]. Recent MEF methods [25, 35, 68, 61, 16, 62, 36] show promise, but still face significant hurdles. Supervised approaches rely on ground-truth (GT) images that are often manually selected or synthetically generated, limiting their ability to generalize to diverse, real-world scenes [31, 30, 29, 28, 37]. Unsupervised methods, on the other hand, using feature-level or pixel-level constraints [1, 3, 73, 74, 75], still suffer from image degradation due to extreme exposures, failing to fully preserve scene information. Overcoming the reliance on ground truth and accurately recovering scene information under extreme exposures remains a critical challenge in MEF.
Figure 1: Traditional Retinex-based techniques for MEF struggle with overexposure, which corrupts the estimated reflectance $R$. Our model addresses this by explicitly modeling glare $G$, thereby restoring a clean shared reflectance $\hat{R}$ for a satisfying result.
@@
The fundamental premise of MEF is that an image sequence captures an invariant scene under varying exposures. This principle aligns naturally with Retinex theory, which postulates that an image $I$ can be decomposed into a product of the illumination component $L$ and the scene reflectance $R$, i.e., $I=L\cdot R$ [11, 23, 26]. This theory is extensively applied in applications such as low-light image enhancement, dynamic range extension, and more [59, 46, 39, 54]. In this context, each $L$ captures the lighting specific to an exposure level, while $R$ represents the scene’s intrinsic, constant properties, which remain consistent under different lighting conditions. Within the Retinex framework, multi-exposure image sequences share a common reflectance component, while the primary distinction among the images is their varying exposure levels. Retinex theory effectively decouples exposure conditions from scene information, eliminating the need for manually designed features and allowing scene details to be revealed without relying on ground truth. Moreover, by isolating the illumination component, the fusion process becomes more flexible and targeted, enabling the generation of images that are not confined to any specific exposure level. Therefore, this paper leverages Retinex theory to address the unsupervised MEF problem, processing $L$ and $R$ distinctly to achieve both faithful scene reconstruction and controllable fusion.
@@
Historically, Retinex theory has been predominantly applied to low-light image enhancement, where the degradation primarily results from reduced illumination levels. However, in MEF, the scope of challenges expands to include glare effects from excessive exposure. These effects elevate pixel values and distort the extraction of the reflectance component $R$, as demonstrated in Fig. 2. This disruption in learning a consistent $R$ leads to artifacts like color shifts and detail loss in Fig. 1. To properly adapt the theory for MEF, our work extends the classic Retinex model to explicitly account for this glare phenomenon.
@@
This paper introduces an unsupervised multi-exposure image fusion method based on Retinex theory, designed to address its key challenges. Our framework models glare effects from excessive exposure, allowing the extracted reflectance component $R$ to align with a shared component $\hat{R}$ instead of converging [12]. To mitigate glare, we employ a bidirectional loss on $\hat{R}$ to preserve the scene’s colors and details. Furthermore, we implement a parameterized illumination fusion criterion that ensures brightness variation and fidelity, facilitating customized exposure adjustments across the image. The contributions of this paper are summarized as follows:
@@
1) We propose an effective Retinex-based unsupervised fusion method that separates reflectance from varying exposure levels, enabling precise exposure blending and improved detail preservation.
2) We introduce a method to mitigate glare in overexposed regions by applying a bidirectional constraint on a shared reflectance, effectively restoring details without pixel-level supervision.
@@
3) We design a flexible fusion strategy using a symmetric and monotonic function to achieve manually controllable exposure adjustments across the image while ensuring contrast is maintained.
4) We conduct extensive experiments on public MEF datasets, demonstrating our method’s superior fusion performance and controllability.
@@
Figure 3: The workflow for the proposed Retinex-MEF. Our method includes three modules: the Illumination Estimation Module for estimating illumination, the Shared Reflection Extraction Module for extracting the common reflectance component of the scene, and the Glare Estimation Module for glare estimation. An image with a random exposure level from multi-exposure images is sampled in each step, enhancing the ability of each module to estimate different components through Retinex-based decomposition and reconstruction.
@@ 2 Related Work
2.1 Deep learning-based MEF
Deep learning-based MEF methods can be categorized into branches of discriminative models [18], generative models [62, 52], and unified fusion models [76, 2]. Discriminative models, built on CNNs [21] or Transformers, learn to either predict fusion weights [18, 44] or directly reconstruct fused images. They can be trained with ground-truth (GT) supervision or in an unsupervised manner [49] using similarity losses like MEF-SSIM [42]. Generative models, such as Generative Adversarial Networks (GANs) [62] and diffusion models [52], generate a fused image whose distribution adversarially driven to be consistent with the source images, ensuring realistic results. Unified fusion models [9, 61, 78] aim to handle multiple fusion tasks in a single framework. They often employ unsupervised strategies, such as source image reconstruction [67, 2], masked image modeling [34], or incorporating guidance from vision language models [76] and task-adaptive losses [2].
@@ 2.2 Retinex theory
Retinex theory has long been a cornerstone of low-light image enhancement. Traditional methods like Single-Scale Retinex (SSR) [20] and Multi-Scale Retinex (MSR) [48] used Gaussian filters to improve contrast but often suffered from artifacts like color distortion. In deep learning frameworks, Retinex serves as a powerful physical prior. Numerous methods leverage neural networks to decompose images into reflectance and illumination [15, 70, 57, 10, 72]. This framework has been integrated with advanced architectures, including Transformers (e.g., Retinexformer [6]), GANs (e.g., EnlightenGAN [19]), and diffusion models (e.g., Diff-Retinex [64]) to tackle low-light enhancement. Specialized models have also emerged to address specific issues like color shifts (CRetinex [63]) or to use deep unfolding techniques (URetinex [59]).
Among these, the most relevant to our work is PairLIE [12], which pioneered unsupervised low-light enhancement by learning a shared reflectance from paired low-light images. Building on this foundation, our work adapts the concept of shared reflectance to the MEF task. We introduce a framework that not only learns a shared reflectance $\hat{R}$ but also explicitly models and mitigates glare. This allows us to restore a clean, consistent reflectance from multi-exposure sequences and achieve controllable exposure fusion, effectively handling both underexposed and overexposed regions.
@@ 2.3 Comparison with Existing Approaches
Compared to existing multi-exposure image fusion methods, our approach does not rely on pixel-level constraints of the source images. Instead, we embrace Retinex theory to adjust exposure levels and restore the scene. By modeling the common reflectance component and the glare phenomenon caused by exposure, we restore the common reflectance component $\hat{R}$ that is unaffected by exposure levels. The adjustable fusion of illumination with $\hat{R}$ leads to superior performance in image fusion. Unlike previous Retinex-based methods, we address the glare phenomenon under extreme lighting conditions and effectively model it. Through bidirectional loss constraints, we extract the common reflectance component $\hat{R}$, providing a more effective Retinex modeling approach for MEF.
@@ 3 Method
While our model is designed to handle multi-level exposure inputs, we focus our discussion on the dual-image case for clarity and comparative purposes. In this setup, in addition to the primary underexposed and overexposed pair $\{I_{u},I_{o}\}\subseteq\{I_{j}\}_{j=1}^{J}$, our method incorporates an auxiliary input, $I_{aux}$. During the training process of our framework, this auxiliary image is randomly sampled from the full exposure sequence $\{I_{j}\}_{j=1}^{J}$ and can therefore be identical to either $I_{u}$ or $I_{o}$. The purpose of this design is to strengthen the model’s ability to form a holistic understanding of the scene. By consistently decomposing and reconstructing $I_{aux}$, our framework learns a more robust estimation of both exposure and reflectance.
@@ 3.1 Overview
Here, $\hat{R}$ symbolizes the intrinsic information of the scene that remains constant, while $G\geq 0$ represents the glare effect induced by exposure. Incorporating $G$ leads to an increase in pixel values in overexposed regions.
@@ 3.2 Retinex-MEF
Estimation of Illumination. Beyond the underexposed and overexposed image pairs $\{I_{u},I_{o}\}$, we have another auxiliary image $I_{aux}\in\{I_{j}\}_{j=1}^{J}$. The exposure level of this auxiliary image is arbitrary and does not require differentiation from $\{I_{u},I_{o}\}$, as its primary function is to facilitate the precise estimation of both illumination and reflectance. The illumination extraction network $\mathcal{N_{I}}(\cdot)$ is composed of five convolutional layers and a $Sigmoid(\cdot)$ activation function, designed to estimate the single-channel illumination $L$ of the auxiliary image $I_{aux}$:
@@ 3.3 Training
According to Eq. (2), the ideal estimation value of $\hat{R}$ is designated as $\tilde{R}$, leading to $I=L\cdot(\tilde{R}+G)$. To align $\hat{R}$ closely with $\tilde{R}$, it is necessary to implement a constraint. Glare effect $G\!\geq\!0$ is not consistent across all exposure levels but is predominant in overexposed images. Thus, at lower exposure levels, $G$ tends towards zero. Consequently, for most exposure levels, $L\cdot\tilde{R}\leq L\cdot(\tilde{R}+G)=I$, though occasionally, $L\cdot\tilde{R}=I$. By applying the constraint $L\cdot\hat{R}\leq I$ universally, we ensure that $\hat{R}\leq\tilde{R}$. To enforce this, we introduce a suppression loss that caps the upper limit of $\hat{R}$, defined as:
@@
The constraint delineated in Eq. (11), which is relevant across all exposure levels, mandates that $\hat{R}\leq\tilde{R}$. Furthermore, in line with the principles of Retinex decomposition, the estimated value of $R$ should be consistent with $\hat{R}$ [59, 64, 12]. Thus, a consistency loss is introduced, assigned a lower weight than $\mathcal{L}_{suppress}$, to ensure adherence to inherent scene consistency mandated by Retinex theory:
@@
This loss ensures that the perturbed reflectance $R$, extracted at varying exposure levels, remains consistent with the shared reflectance $\hat{R}$. Given that $\tilde{R}<R$, the objective of this loss is to maximize $\hat{R}$ within the bounds set by $\tilde{R}$. The constraints $\mathcal{L}_{suppress}$ and $\mathcal{L}_{consist}$ synergistically regulate $\hat{R}$, ensuring it closely approximates $\tilde{R}$. On one hand, with both constraints using the $\ell_{1}$-norm but $\mathcal{L}_{consist}$ assigned a smaller weight than $\mathcal{L}_{suppress}$, $\mathcal{L}_{suppress}$ predominates in conflicts, maintaining $\hat{R}\leq\tilde{R}$. Conversely, when $\hat{R}\leq\tilde{R}$ has been achieved and considering $\hat{R}\leq R$, $\mathcal{L}_{consist}$ acts to elevate $\hat{R}$, guiding it towards convergence with $\tilde{R}$. This synergistic effect is illustrated in Fig. 4.
Figure 4: A schematic illustration of the synergistic effect between suppression loss $\mathcal{L}_{suppress}$ and consistency loss $\mathcal{L}_{consist}$: When $\hat{R}$ exceeds $\tilde{R}$, the two losses conflict, but $\mathcal{L}_{suppress}$ plays a dominant role. Conversely, when $\hat{R}$ is less than $\tilde{R}$, $\mathcal{L}_{consist}$ ensures that $\hat{R}$ increases.
@@
where $\alpha_{1},\alpha_{2},\alpha_{3},\alpha_{4}$ are weighting coefficients $(\alpha_{3}\!>\!\alpha_{4})$.
@@
MEFB Multi-Exposure Image Fusion Dataset [69] SICE Multi-Exposure Image Fusion Dataset [5]